# Automated Bug Generation in the era of Large Language Models

### v1

@@ 5.4. RQ3: Effectiveness in Generating Hard-to-Repair Bugs
[enhanced, frame hidden, boxsep=0pt]Summary. When applied to the same method, BugFarm generates bugs that are harder to repair by learning-based repair techniques compared to alternative approaches. The power of BugFarm will become more evident when the methods are longer, letting it change multiple locations in the method to introduce the bug.
5.5. RQ4: Necessity of Prompt Engineering
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decoratedBag().add(transformer.transform(value));
    return dec;
(a) Incorrect method call
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final Object value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(b) Incorrect loop variable type
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        // Removed the array initialization
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(c) Removed array initialization
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            // Removed the method call within the loop
    return dec;
(d) Removed method call
public static <E> List<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(e) Incorrect return type
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final Object value : values) {
            dec.decoratedBag().add(transformer.transform(value));
    return dec;
(f) Combination of bug 2(a) and 2(b)
Figure 3. The bugs generated without prompt engineering for the code in Figure 0(a).
The quality of prompts significantly impacts the quality of the model’s response (Liu et al., 2023); the more context you provide to the model in the prompt, the better the response it produces. In the context of our problem, identifying the location of Least Attended Statements (LASs) brings three benefits: (1) generating bugs with a similar representation compared to the original code (hard to detect bug), (2) enforcing the model to change multiple lines (hard to repair bug), and (3) reducing the search space, i.e., the location in which the bug can be injected. This will help the model focus on the most important part of the code and generate fewer yet more effective bugs.
Our previous research questions confirm the ability of prompt engineering to generate complex bugs. Generating those bugs without prompt engineering will require more time and cost more. For example, consider prompting an LLM to generate a bug for the code illustrated in Figure 0(a), implementing a logic through $13$ lines of code (some lines are split for better presentation). There are two least attended statements (LASs) in this method, highlighted in green. The probability that LLM changes one of them is $\frac{2}{13}=0.15$, i.e., with a probability of $15\%$, LLM can generate a hard-to-detect bug. However, to ensure the bug is also hard to repair, LLM should ideally modify the two highlighted lines. The probability of LLM choosing the two statements highlighted in green (the LASs) is only 1%, $\frac{1}{\binom{13}{2}}=0.01$. The code snippets in Figure 3 show different bugs generated by prompting GPT-3.5-turbo six times (buggy lines highlighted in red). We used a simple, not crafted prompt, asking it to generate a compilable bug from the given method. We can see that the LLM is quite creative in injecting different bugs (statement deletion, changing types, and incorrect method calls) into the code. But none of the responses change any of the least attended statements and, hence, have no overlap with the BugFarm bugs.
In general, for a given method with $n$ statements, the bug-inducing change can be applied to any line. BugFarm selects only $k$ percent of statements to be changed and instructs LLM to modify them, reducing the search space greatly. With the threshold value $k=20\%$, we can reduce the search space by $80\%$. Without attention analysis, a naive approach for search space reduction would randomly select $k$ percent of the statements. If not among the least attended statements, they can be easily detected or repaired by learning-based techniques, as shown in previous research questions.
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Prompt crafting help BugFarm in generating hard-to-detect and hard-to-repair bugs. The chance of generating a similar dataset without prompt engineering is very narrow, if not impossible, due to the relatively large search space for moderate to long methods.
Figure 4. Performance of BugFarm compared to alternative approaches in unconfirmed bug generation
5.6. RQ5: Performance
To assess the performance characteristics of BugFarm, we measured the time required to extract attention weights from the models and the time it takes to prompt the LLM (GPT-3.5-turbo). We also compared the total time for bug generation in BugFarm compared to alternative approaches. All the experiments were performed on a workstation with NVIDIA GeForce RTX $3090$ GPUs ($24$GB GDDR6X memory) and $24$ 3.50GHz Intel 10920X CPUs ($128$ GB of memory). Figure 4 shows the results, where the red dashed line indicates the mean value. Looking at Attention Analyzer in Figure 4, we can see that it takes $68$, $69$, and $86$ milliseconds on average for BugFarm to extract and analyze attention weights from CodeBERT, CodeT5, and NatGen models. Prompting the LLM for every method also takes about $9$ seconds on average (Prompting box chart at the middle). The prompting time can be affected by multiple factors, including the traffic on the model and prompt size, i.e., the number of tokens in the prompt.
We also measured the total time each technique takes to generate unconfirmed bugs per method (End-to-End Unconfirmed Bug Generation in Figure 4). Compared to attention analysis and prompting, the overhead of other sub-components in BugFarm, such as parsing and bug selection, is negligible. Therefore, the average time for generating all the BugFarm unconfirmed bugs per each method is $9.29$ seconds (mostly dominated by the prompting time). In comparison, it takes LEAM and $\mu$BERT $35$ and $28$ seconds on average to generate unconfirmed bugs per method. This is because these approaches generate more unconfirmed bugs, as illustrated in Table 1 under #UB columns. For LEAM, this time does not include the training time of the model, which is $24$ hours  [^8] This number is quoted from the paper.. Also, since $\mu$BERT takes a long time to mutate big classes, we put a timeout of $15$ minutes on it to avoid a long generation time. Validation of unconfirmed bugs was also very time-consuming; we spent around $80,000$ CPU core-hours to validate over $2.5$ million mutants from $\mu$BERT, LEAM, and BugFarm.
{tcolorbox}
[enhanced, frame hidden, boxsep=0pt]Summary. Compared to alternative approaches, BugFarm is an efficient and scalable technique for generating bugs. More precisely, BugFarm is 74% and 67% faster in end-to-end unconfirmed bug generation than LEAM and $\mu$BERT, respectively.
5.7. Discussion
Our comprehensive evaluation of BugFarm compared to the two most recent alternative approaches confirms the ability of generated bugs to challenge learning-based bug prediction and repair techniques. We want to emphasize that our objective in this paper is to generate hard to detect and repair bugs, concerning learning-based techniques. As a result, comparing with real-world bugs is out of the scope of this paper. We have no claim that BugFarm bugs mimic real-world bugs (because they do not need to) or that they do not. In fact, BugFarm leverages GPT-3.5-turbo for generating bugs, which theoretically have seen many real-world bugs in its training data. This can potentially help BugFarm bugs be similar to real-world bugs compared to $\mu$BERT, which does not concern such similarity. We have identified all the potential threats to the validity in the design of the evaluation and have addressed them throughout the evaluation section when it was the most relevant. Our next step would be investigating the extent to which BugFarm bugs can enhance downstream applications of synthetic bugs in learning-based techniques, such as test case prioritization and bug localization.
6. Related Work
Our research is related to prior work on (1) real-world bug/vulnerability benchmarks, (2) mutation testing, and (3) learning-based bug/vulnerability generation.
Real-world bug/vulnerability benchmarks. Several attempts have been made to construct real-world bug datasets manually. Defects4J (Just et al., 2014a), BugSwarm (Tomassi et al., 2019), Bugs.jar (Saha et al., 2018), and RegMiner (Song et al., 2022) are the commonly used Java bug datasets that have mined GitHub to collect regression bugs from bug-fixing commits. BigVul (Fan et al., 2020) and CVEFixes (Bhandari et al., 2021) are real-world examples of vulnerabilities collected from bug-fixing reports in the CVE/NVD database (Booth et al., 2013). BugFarm complements the bugs and vulnerabilities in these datasets with hard-to-repair and hard-to-detect bugs. Also, the bugs in these datasets only represent human mistakes, which could be potentially different from the mistakes AI programming tools make.
Mutation testing. Mutation testing has been widely used in testing programs written in different languages (Deng et al., 2017; Schuler and Zeller, 2009; Just, 2014; Coles et al., 2016; Delamaro et al., 2001; Ma et al., 2005; Mirshokraie et al., 2013; Deng et al., 2015; Black, 2017; Okun et al., 2013), as well as testing program properties such as specifications (Martin and Xie, 2007), memory usage (Wu et al., 2017), and energy consumption (Jabbarvand and Malek, 2017). Syntactic mutation operators, where small syntactic changes in the code generate the artificial bugs, are not representative of real bugs (Just et al., 2014b). Constructing defect-based mutation operators alleviates the mentioned limitation but requires domain knowledge. If done manually, it also can be labor-intensive and time-consuming. BugFarm is an automated technique for generating bugs; hence, it is superior to non-automated mutation testing. The current implementation of BugFarm does not generate specific categories of bugs. However, by incorporating some domain knowledge as additional contexts to our prompts, one can expand the bug injection component of BugFarm to generate specific categories of bugs.
Learning-based bug/vulnerability generation. Learning-based bug generation techniques were first proposed to overcome the limitations of manual defect model construction in mutation testing (Brown et al., 2017). Such techniques leverage machine learning to learn the bug or vulnerability patterns from real-world bug fixes and generate mutants accordingly. DeepMutation (Tufano et al., 2019) is such a technique that relies on sequence-to-sequence neural machine translation for learning and generating bugs. SemSeed (Patra and Pradel, 2021) is a technique that extracts bug patterns from real-world bug fixes and injects them into other programs so that the bug in the new program is syntactically different but semantically similar. MutationMonkey (Beller et al., 2021) mines bug patterns from historical changes and transforms them into mutation operators semi-automatically.
VulGen (Nong et al., 2023) combines pattern mining and deep learning to generate realistic bugs. Specifically, it mines a large corpus of vulnerability-fixing commits to extract bug patterns, and trains a deep learning model on the same dataset to learn where to inject the bugs. LEAM (Tian et al., 2022) learns to mutate code from large examples of real-world bug-fixing commits. To generate syntactically correct bugs, LEAM represents code as a sequence of AST nodes and learns to apply grammar rules to select and modify the code. Finally, $\mu$BERT (Khanfir et al., 2023) produces buggy versions by replacing code tokens with the spacial <mask> token, and uses CodeBERT to predict the masked token. Both LEAM and $\mu$BERT incorporate beam search (Freitag and Al-Onaizan, 2017) to generate bugs that involve more than one statement.
@@ 6. Related Work
7. Concluding Remarks
Bug benchmarks are essential in software engineering to evaluate automated techniques concerning bugs. The advent of learning-based software engineering demands even more for automated bug-generation techniques, not only for evaluation but also for training or fine-tuning ML4SE models. In this paper, we presented BugFarm, a model-in-the-loop technique for the automated generation of hard-to-detect and hard-to-repair bugs. Our empirical evaluation shows the superiority of BugFarm compared to alternative bug generation approaches in (1) generating unique and high-quality bugs, (2) constructing bug benchmarks that are complex both in terms of detection and repair, and (3) injecting hard-to-detect bugs into arbitrary code. BugFarm does not rely on existing bug datasets and is model- and programming language-agnostic.
We believe our novel perspective on the bug generation problem offers several directions for future work. First of all, the focus of this paper is on attention analysis to identify the weak spots of transformer-based models. While state-of-the-art code-language models and LLMs are all based on transformers, we currently consider alternative approaches generalized to other neural architectures. Second, we plan to evaluate the quality of our bugs for improving other software engineering tasks, including bug localization, test oracles, and test-suite management. Based on the promising results in bug detection and repair, we are confident that our generated bugs are useful for better evaluation and challenging those tasks.

### v2

@@ V-D RQ3: Effectiveness in Generating Hard-to-Repair Bugs
Summary. When applied to the same method, BugFarm generates bugs that are harder to repair by learning-based repair techniques compared to alternative approaches. The power of BugFarm will become more evident when the methods are longer, letting it change multiple locations in the method to introduce the bug.
V-E RQ4: Performance
We measured the time required to extract attention weights from the models and the time it takes to prompt the LLM (GPT-3.5-turbo). We also compared the total time for bug generation in BugFarm compared to alternative approaches. All the experiments were performed on a workstation with NVIDIA GeForce RTX $3090$ GPUs ($24$GB GDDR6X memory) and $24$ 3.50GHz Intel 10920X CPUs ($128$ GB of memory). Figure 4 shows the results, where the red dashed line indicates the mean value. It takes $68$, $69$, and $86$ milliseconds on average for BugFarm to extract and analyze attention weights from CodeBERT, CodeT5, and NatGen models. Prompting the LLM for every method also takes about $9$ seconds on average (Prompting box chart at the middle). The prompting time can be affected by multiple factors, including the traffic on the model and prompt size, i.e., the number of tokens in the prompt.
We also measured the total time each technique takes to generate mutants (unconfirmed bugs) per method (End-to-End Mutant Generation in Figure 4). Compared to attention analysis and prompting, the overhead of other sub-components in BugFarm, such as parsing and bug selection, is negligible. Therefore, the average time for generating all the BugFarm mutants per each method is $9.29$ seconds (mostly dominated by the prompting time). In comparison, it takes LEAM and $\mu$BERT $35$ and $28$ seconds on average to generate mutants per method. This is because these approaches generate more mutants, as illustrated in Table I under #SCM columns. For LEAM, this time does not include the training time of the model, which is $24$ hours  [^9] This number is quoted from their paper.. Also, since $\mu$BERT takes a long time to mutate big classes, we put a timeout of $15$ minutes on it to avoid a long generation time. Automated validation was also very time-consuming; we spent around $60,000$ CPU core-hours to validate over $1.9$ million mutants from $\mu$BERT, LEAM, and BugFarm.
Summary. Compared to alternative approaches, BugFarm is an efficient and scalable technique for generating bugs. More precisely, BugFarm is 74% and 67% faster in end-to-end mutants generation than LEAM and $\mu$BERT, respectively.
Figure 4: Performance of BugFarm compared to alternative approaches in mutant generation
V-F Discussion
BugFarm aims to generate hard-to-detect and repair bugs concerning learning-based techniques. Our bugs are not designed to replace real-world bugs but complement them from a new perspective, i.e., having a very close code representation as the original code while different in several places. We have no claim that BugFarm bugs mimic real-world bugs (because they do not need to). As a result, comparing with real-world bugs is out of the scope of this paper. However, BugFarm leverages GPT-3.5-turbo for generating bugs, which theoretically have seen many real-world bugs during training. This can potentially help BugFarm bugs be similar to real-world bugs compared to $\mu$BERT, which does not concern such similarity.
VI Related Work
Real-world bug/vulnerability benchmarks. Several attempts have been made to construct real-world bug datasets manually. Defects4J [23], BugSwarm [40], Bugs.jar [34], and RegMiner [37] are the commonly used Java bug datasets that have mined GitHub to collect regression bugs from bug-fixing commits. BigVul [14] and CVEFixes [6] are real-world examples of vulnerabilities collected from bug-fixing reports in the CVE/NVD database [8]. BugFarm complements the bugs and vulnerabilities in these datasets with hard-to-repair and hard-to-detect bugs. Also, the bugs in these datasets only represent human mistakes, which could be potentially different from the mistakes AI programming tools make.
Learning-based bug/vulnerability generation. Learning-based bug generation was proposed to overcome the limitations of manual defect model construction [9]. Such techniques learn the bug or vulnerability patterns from real-world bug fixes and generate mutants accordingly. DeepMutation [41] is a technique that relies on sequence-to-sequence neural machine translation for learning and generating bugs. SemSeed [31] extracts bug patterns from real-world bug fixes and injects them into other programs so that the bug in the new program is syntactically different but semantically similar. MutationMonkey [5] mines bug patterns from historical changes and transforms them into mutation operators semi-automatically. VulGen [30] combines pattern mining and deep learning to generate realistic bugs. LEAM [39] learns to mutate code from large examples of real-world bug-fixing commits. $\mu$BERT [24] produces buggy versions by replacing code tokens with the spacial <mask> token, and uses CodeBERT to predict the masked token. Both LEAM and $\mu$BERT incorporate beam search [16] to generate bugs that involve more than one statement.
@@ VI Related Work
VII Concluding Remarks
Bug benchmarks are essential in software engineering to evaluate automated techniques concerning bugs. The advent of learning-based bug-related techniques demands automated bug-generation techniques for proper evaluation. In this paper, we presented BugFarm, a model-in-the-loop technique for the automated generation of hard-to-detect and repair bugs. Our empirical evaluation shows the superiority of BugFarm to alternative mutant generation approaches in generating unique and high-quality hard-to-detect and repair bugs. BugFarm does not rely on existing bug datasets and is model- and programming language-agnostic.